Find the value of x.
logx4/169=2
a) x=3/13
b) x=2/13
c) x=+-13/2
d) x=13/2
e) x=+-14/2

Answer by stanbon(48569)

Find the value of x.
logx4/169=2
---
x^2 = (4/169)
x = 2/13
==============
Cheers,
Stan H.